OpenStudy (anonymous):

What is the equation of the hyperbola with a=6 and c=9? assume that the transverse axis is horizontal

OpenStudy (anonymous):

c^2-a^2=b^2 b^2=81-36=45 x^2/36-y^2/45=1?